Let an-nin 2 1. Evaluate the following 4 an +WE n=1
Answer #1

welcome

given

an = (n -1) / (n + 1)

thus

a1 = (1-1) / (1+1) = 0 / 2 = 0

a2 = (2 - 1) / ( 2+ 1) = 1 / 3

a3 = (3 - 1) / (3 + 1) = 2 / 4

a4 = ( 4-1) / ( 4+1) = 3 / 5

thus

a1 + a2 + a3 + a4 = 0 + (1/3) + ( 2/4) + (3/5)

= ( 20 + 15 + 12) / 60

= 47 / 60